OpenStudy (anonymous):

A box that is to be mailed at the standard rate must be less than 12 inches on each edge. Which of the following expressions represents the volume of the box? V < 1 ft3 V > 1 ft3 V ≤ 1 ft3 V < 12 in3

OpenStudy (kropot72):

What is the volume of a box in the shape of a cube with each edge 1 ft long?

OpenStudy (anonymous):

volume is length cubed so \[V<L ^{3}\] hence \[V<12^{3}inches\] dont know what a foot is in inches as i use metric system but i assume answer is D

OpenStudy (kropot72):

\[V<12^{3}inches\]is not one of the choices. Don't confuse it with the choice V < 12 in3 which stands for 12 cubic inches.

OpenStudy (anonymous):

ok i dont get imperial system join modern age and use metric it makes much more sense

OpenStudy (kropot72):

If a box in the shape of a cube has each side equal to 12 inches the volume is: 12 * 12 * 12 = 1728 cubic inches.

OpenStudy (anonymous):

yeah that what i meant but what is a cubic foot

OpenStudy (kropot72):

One cubic foot has a volume equal to that of a cube with each side measuring 1 foot. 1 * 1 * 1 = 1 cubic foot. There are 12 inches in a foot. Therefore 1 cubic foot = 12 * 12 * 12 = 728 cubic inches.